OpenStudy (anonymous):

find the coordinate of vertex for the parabola defined by the given quad. function. F(x)=-x^2-8x+2

OpenStudy (anonymous):

is the answer (-4,18)

OpenStudy (anonymous):

\[- \frac{ b }{ 2a } =-\frac{ -8 }{ 2(-1) }\]

OpenStudy (anonymous):

answer is correct:)

OpenStudy (anonymous):

so -(-8)/-1=-8 then divided -8/2 =-4

OpenStudy (anonymous):

thank you!